OpenStudy (anonymous):

If the revenue of a shoe company is f(x) = 33x + 16, where x is the number of pairs of shoes sold, and the expenses of the company are g(x) = 4x + 50, where x is also the number of pairs of shoes sold, what is the profit of the company if they sell 105 pairs of shoes? A. $3,481 B. $3,011 C. $3,951 D. $3,111

OpenStudy (aaronq):

you wanna subtract the expenses from the profit so, find f(x)-g(x) when x=105

OpenStudy (anonymous):

huh?

OpenStudy (aaronq):

f(x) = 33x + 16; (x) = 4x + 50 profit=f(x)-g(x)= 33x + 16 - (4x + 50) =33x + 16 - 4x - 50

OpenStudy (anonymous):

=29x-44?

OpenStudy (aaronq):

yeah, now find what it is when x=105 profit=29(105)-44

OpenStudy (aaronq):

it's 29x-34, not -44

OpenStudy (anonymous):

lol. okay. Then divide both sides by 29?

OpenStudy (aaronq):

you're not solving for x, just multiply and subtract profit = 29(105)-34